OpenStudy (anonymous):

a store mixes kenyan coffee worth $11 per kilogram nd turkish coffee worth $14 per kilogram. The mixture is to sell for $12 per kilogram. Find how much of each should be used to make a 228 kilogram mixture

OpenStudy (beginnersmind):

How much Kenyan coffee and how much Turkish coffee should go into 1 kg of mixture?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

we have to introduce a variable. lets call the number of kilos of turkish coffee \(x\) so then the number of kilos of kenyan must be \(228-x\) becauseteh total is 228

OpenStudy (anonymous):

if you have \(x\) kilos of turkish, then it cose \(14\times x\) and the \(228-x\) kilos of kenyan cost \(11\times (228-x)\)

OpenStudy (anonymous):

all together you want the total cost to be \(12\times 228=2736\) so set \[14x+11(228-x)=2736\] and solve for \(x\)